OpenStudy (anonymous):

Calculate the amount of heat needed to raise the temperature of 358g of ice from -24.5degreesC to 456degreesC.

OpenStudy (greatlife44):

@Clemsongator Do you know how to approach this?

OpenStudy (anonymous):

i got the amount of energy to cget it to 0C as 18419.1J but idk where to go from there

OpenStudy (greatlife44):

alright. The amount of heat released q is equal to the mass of the substance in grams, times C the specific heat times delta T the temperature change. \[q = m~C~\Delta~T\]

OpenStudy (anonymous):

yeah

OpenStudy (greatlife44):

Our final temperature is 45.6 degrees C right?

OpenStudy (anonymous):

456 degrees C

OpenStudy (greatlife44):

Well, I guess I decided to find the specific heat of ice. how much energy we need to put into ice to melt it. we can't do our calculation without it. 458*C is a pretty high temperature so that's why i'm asking. \[2.65~ J/grams*K\] specific heat of ice K means temperature in kelvins. so we convert our temperatures to kelvins, \[C^{0}+273 = Kelvins\] \[-24.5+273 = 248.5 K_{T_{1}}\] \[273+456 = 729K_{T_{2}}\] then we put everything into this formula alright. \[q = m~C~\Delta~T\] \[q = (358g)(2.65 \frac{ J }{ grams*K })*(729_{t2}-248.5_{t1})K = 456~Kilo~Joules\]

OpenStudy (anonymous):

i thought the specific heat of ice was 2.1

OpenStudy (greatlife44):

The values are different depending on what units you choose to use @Clemsongator